Three apples cost as much as 4 pears. Three pears cost as much as 2 oranges. How many apples cost as much as 72 oranges?
I think it is 36. Am I correct? Thanks!

Answer by macston(5194)

1 pear=3/4 apple
1 orange=2/3 pear
1 orange=(2/3)(3/4 apple)=1/2 apple
So 72 oranges cost as much as 36 apples.